OpenStudy (anonymous):

Twelve members of the marching band are on the football field. Eight of the band members are flutists and four are drummers. IN how many ways could you select a group consisting of exactly three flutists and two drummers?

OpenStudy (anonymous):

I don't think that's right Maph It is (8 choose 3) times (4 choose 2) which is 336

OpenStudy (anonymous):

\[\left(\begin{matrix}8 \\ 3\end{matrix}\right) = 8!/(3!5!) * 4!/(2!2!) = 336\]